Price of Computers Price of Speakers Price of Monitors 2006 $1200 $70 $150 2011 800 S40 $100 Table-9 ble-9: Suppose we have a consumer basket with 1 computer, 3 speakers, and 2 monitors. Compute the CPI using 2006 as a base year Irom Table-9 A 67 C 151 D 65
Answer #1

Image result for cpi formula

In the base year (2006)

Cost of consumer basket = price of computer * 1 + price of speaker * 3 +price of monitor * 2

= 1200*1 + 70*3 + 150*2 = 1710

In the given year (2011)

Cost of consumer basket = price of computer * 1 + price of speaker * 3 +price of monitor * 2

= 800*1 + 40*3 + 100*2 = 1120

710 100 152.63 1120 * 100 152.68 CPI =